A 32.9-kg child swings on a rope with a length of 6.11 m that is hanging from a tree. At the bottom of the swing, the child is moving at a speed of 4.2 m/s. What is the tension in the rope?
Answer #1

T=mv^2/r=centripetal force+mg

=32.9*4.2^2/6.11=94.98 +32.9*9.8=416.8 N
Answer #2

T=\frac{m.v^2}{r}

m=32.9

r=6.11 m

v= 4.2 m/sec

T=32.9x(4.2)^2/6.11

=94.98 Newton